OpenStudy (anonymous):

What are the binomial factors of the expression below? x^2 - 2x - 15 A.) (x-3)(x-5) B.) (x+3)(x+5) C.) (x-3)(x+5) D.) (x+3)(x-5) Help? Best answer will receive a medal.

OpenStudy (anonymous):

x2 + 2x = 15 This is a quadratic equation which can be solved 3 ways - factoring, graphing, quadratic formula. I will use factoring. We need the equation to equal 0 to solve it, so let's subtract 15 from each side: x2 + 2x - 15 = 0 This is a trinomial (3 terms) which we want to factor into 2 binomials (2 terms). It will look like (x + )(x - ) = 0 but we need to fill in the blanks with factors of 15. We have + then - because we want -15 when we multiply the 2 factors and the only way to get a negative is to multiply a positive times a negative. (3)(5) = 15 so let's try that. (x + 3)(x - 5) = x2 - 5x + 3x - 15 = x2 - 2x - 15 close but not correct We need +2x in the middle, so let's try swapping the 3 and the 5 (x + 5)(x - 3) = x2 - 3x + 5x - 15 = x2 + 2x - 15 correct! Now let's solve: for (x + 5)(x - 3) = 0, either x + 5 =0 or x - 3 = 0 If x + 5=0, then solve for x and get x = -5 If x - 3 = 0 then solve for x and get x = 3 So your complete solution is x = -5 and x = 3

OpenStudy (anonymous):

Thanks!